# Polish Medical Exams: A new dataset for cross-lingual medical knowledge transfer assessment

### v1

@@ 7 Cross-lingual knowledge transfer
To compare the performance of various LLMs on Polish and English versions of the same datasets, we restricted the LEK and LDEK datasets to identical subsets (LEK exams in English are exact translations of the Polish exams). The analysis results, similar to the previous one, are presented in Tables 6 and 7. As shown, all medical models, except for OpenMeditron/Meditron3-70B, perform better on the English versions of the datasets. This may be due to these models being fine-tuned on English medical corpora. General-purpose multilingual models perform better on the English versions of the exams as well. This result is anticipated since these models are trained on corpora containing significantly more English than Polish. While these models are proficient in Polish, their performance on the tests remains lower in Polish than in English. The difference can be considerable; for example, meta-llama-Meta-Llama-3.1-8B-Instruct passed only one LEK exam in Polish but passed all 13 when translated into English. However, as model quality improves, the performance gap between languages narrows. For instance, with meta-llama-Meta-Llama-3.1-8B-Instruct, the accuracy difference between Polish LEK (51.25%) and English LEK (64.69%) is 13.44 percentage points (or a 26% relative change). In contrast, with meta-llama-Meta-Llama-3.1-70B-Instruct, the difference is only 1.66 percentage points (80.94% for Polish LEK vs. 82.60% for English LEK, or a 2% relative change).
For GPT-4o-mini, which generally performs well, the results in English are only slightly better than in Polish. Interestingly, for GPT-4o, performance is actually higher on the Polish version. The only Polish LLM, Bielik, performs better on Polish LEK and slightly better on Polish LDEK, likely due to its fine-tuning from the multilingual model Mistral-7B-v0.2 specifically for Polish. Overall, our observations suggest that language transfer is more effective as the model’s general performance improves.
@@
meta/llama-Meta-Llama-3.1-70B-Instruct 80.94 82.60 61.75 63.17
meta/llama-Meta-Llama-3.1-8B-Instruct 51.25 64.69 41.06 47.71
@@
Table 6: The comparison of LLMs on Polish and English datasets, using the same LEK and LDEK exams, is represented as a percentage of correct answers. The English versions of these exams are translated from the Polish originals by human medical experts, specifically the exam creators. The language version of the exam on which each model performs better is bolded.
@@
meta-llama-Meta-Llama-3.1-70B-Instruct 13/13 13/13 12/13 13/13
meta-llama-Meta-Llama-3.1-8B-Instruct 1/13 13/13 0/13 0/13
@@
Table 7: The comparison of LLMs on Polish and English datasets using the same LEK and LDEK exams is represented as a passed exams. The English versions of these exams are translated from the Polish originals by human medical experts, specifically the exam creators. The language version of the exam on which each model performs better is bolded.
@@ 8 Comparison against human results
For each group of models described in Section 5, the top-performing model was selected to compare its predictions against the exam results of human students from the last four LEK and LDEK sessions: Spring 2024, Autumn 2023, Spring 2023, and Autumn 2022. That is 977 LEK and 984 LDEK questions. The selected models are as follows: Meditron3-70B for medical-specific models, Meta-Llama-3.1-70B-Instruct for general-purpose multilingual models, Bielik-11B-v2.2-Instruct for Polish-specific models, and gpt-4o-2024-08-06 for restricted API models.
@@
While all selected models passed the chosen LEK exams, only Meta-Llama-3.1-70B-Instruct and gpt-4o-2024-08-06 scored within the range defined by an average number of points $\pm$ standard deviation achieved by humans. Assuming a normal distribution of exam results, it could be concluded that these models performed as a typical medical student. Notably, for the spring 2024 LEK exam, Meditron3-70B also achieved an average-level result, while gpt-4o-2024-08-06 exceeded the average student score. These findings are presented in Table 9.
For the LDEK exams, all models performed noticeably worse. Assuming a normal distribution of exam results, only gpt-4o-2024-08-06 maintained a performance level comparable to that of an average medical student, consistent with its LEK exam results. In contrast, Meditron3-70B and Bielik-11B-v2.2-Instruct performed poorly, failing all exams, while Meta-Llama-3.1-70B-Instruct scored below the average but managed to pass each exam. These outcomes are summarized in Table 10.
@@
The same models were used to compare their performance with that of human students on the PES exams. This analysis was conducted on a dataset created from the intersection of human results and LLM test results, encompassing 9,965 medical questions across 68 specializations from 12 exam sessions: Spring 2024, Autumn 2023, Spring 2023, Autumn 2020, Autumn 2019, Spring 2018, Spring 2017, Autumn 2016, Spring 2016, Autumn 2015, Spring 2012, and Autumn 2008. The number of specializations is smaller than in the previous analysis due to inconsistencies in the specialization names between shared exams and published human results. The best-performing model was gpt-4o-2024-08-06, which achieved results in 60% of cases better than half of the student population or within the top 25% of scores. Notably, this model outperformed all students in a thoracic surgery exam. However, it is important to note that the student population for this particular exam was relatively small, consisting of only six participants. However, it is worth noting that even the best model achieved results worse than half of the student population in over 30% of specializations. For the Audiology & phoniatrics specialization, the model underperformed compared to all students. However, the student population for that particular case was relatively small, consisting of only nine participants. The second-best model, Meta-Llama-3.1-70B-Instruct, delivered significantly worse performance compared to the best model. Only 10% of its results across specializations were above the population median, while in over 30% of medical specializations, its performance was above the 25th percentile. The remaining models, Meditron3-70B and Bielik-11B-v2.2-Instruct, performed extremely poorly, with most of their results falling below the 25th percentile or even below the lowest scores of the entire student population. The students’ results are illustrated in the box plot diagrams presented for each medical specialization in Appendix C. In these plots, the whiskers represent the minimum and maximum student scores rather than the inter-quartile range. The red horizontal line indicates the exam passing threshold, while the other colored horizontal lines represent the scores achieved by the tested LLM models. An aggregated exam results are provided in Table 8, where the scores of each model $X$ are categorized into the following ranges:
  • $X<p_{1}$: Represents the range where model $X$ achieves scores lower than all students, indicating that the model underperforms compared to humans.
  • $X\in[p_{0},p_{25})$: Results achieved by model $X$ fall within the range of the lowest 25% of student scores.
  • $X\in[p_{25},p_{50})$: Results achieved by model $X$ are between the 25th and 50th percentiles of student scores, indicating performance worse than half of the student population but better than the first quartile.
  • $X\in[p_{50},p_{75})$: Results achieved by model $X$ fall in the range of scores better than half of the students but below the top 25%.
  • $X\in[p_{75},p_{100})$: Results achieved by model $X$ are within the top 25% of student scores.
  • $X\geq p_{100}$: Results where model $X$ achieved or exceeded the best human score.
@@
OpenMeditron/Meditron3-70B $X<p_{1}$ 16 23.53%
$X\in[p_{0},p_{25})$ 48 70.59%
$X\in[p_{25},p_{50})$ 2 2.94%
$X\in[p_{50},p_{75})$ 2 2.94%
$X\in[p_{75},p_{100})$ 0 0%
$X\geq p_{100}$ 0 0%
meta-llama/Meta-Llama-3.1-70B-Instruct $X<p_{1}$ 5 7.35%
$X\in[p_{0},p_{25})$ 33 48.53%
$X\in[p_{25},p_{50})$ 23 33.82%
$X\in[p_{50},p_{75})$ 5 7.35%
$X\in[p_{75},p_{100})$ 2 2.94%
$X\geq p_{100}$ 0 0%
speakleash/Bielik-11B-v2.2-Instruct $X<p_{1}$ 24 35.29%
$X\in[p_{0},p_{25})$ 43 63.24%
$X\in[p_{25},p_{50})$ 1 1.47%
$X\in[p_{50},p_{75})$ 0 0%
$X\in[p_{75},p_{100}])$ 0 0%
$X\geq p_{100}$ 0 0%
gpt-4o-2024-08-06 $X<p_{1}$ 1 1.47%
$X\in[p_{0},p_{25})$ 11 16.18%
$X\in[p_{25},p_{50})$ 13 19.12%
$X\in[p_{50},p_{75})$ 21 30.88%
$X\in[p_{75},p_{100})$ 21 30.88%
$X\geq p_{100}$ 1 1.47%
Table 8: Aggregated exam results categorizing model $X$ performance relative to the student population across various percentiles, from scores below all students ($X<p_{1}$) to scores compared to or exceeding the best human results ($X\geq p_{100}$).
@@
with standard deviation ±$\pm 19.79$$\pm 18.38$$\pm 18.66$$\pm 16.95$
Table 9: Comparison of top-performing LLMs and average human results, including standard deviation, across four selected LEK exams. Red represents values below the passing threshold of 112 points, orange highlights scores below average minus one standard deviation, green indicates scores above average plus one standard deviation, and black represents scores within one standard deviation of the average.
@@
with standard deviation ±$\pm 26.08$$\pm 19.08$$\pm 21.13$$\pm 23.52$
Table 10: Comparison of top-performing LLMs and average human results, including standard deviation, across four selected LDEK exams. Red represents values below the passing threshold of 112 points, orange highlights scores below average minus one standard deviation, green indicates scores above average plus one standard deviation, and black represents scores within one standard deviation of the average.
@@
While LLMs have demonstrated impressive performance on Polish medical multiple-choice exams, this achievement represents only a narrow facet of medical expertise. Becoming a licensed physician in Poland requires extensive training, rigorous coursework, and hands-on experience with practical medical procedures—far beyond what written exams can assess. Clinical practice necessitates analyzing diverse information and solving complex problems with multiple possible solutions. Physicians must determine what data is needed, obtain it through patient interviews, physical examinations, diagnostic tests, and consultations—all heavily reliant on direct human interaction that AI models cannot replicate. Moreover, the exams are multiple-choice, and real-world work is not narrowed to a few possible options. Therefore, despite strong exam results, LLMs cannot currently substitute the comprehensive qualifications and essential human interactions integral to effective medical care. However, this work shows that LLMs may be useful tools for medical practitioners. Ullah et al. (2024); Park et al. (2024); Clark and Bailey (2024); Liu et al. (2023); Lee et al. (2023).
@@
Ethics Statement
This research involves the collection and analysis of publicly available medical examination data from the Medical Examination Center and the Supreme Medical Chamber. All data used are openly accessible and do not contain any personally identifiable information or confidential content. We have ensured compliance with the terms of use specified by the data providers and have respected all applicable laws and regulations regarding data privacy and intellectual property.
The primary objective of this study is to assess the performance of LLMs on Polish medical exams and to introduce a new dataset for cross-lingual knowledge transfer assessment. While LLMs have demonstrated strong performance on multiple-choice medical exams, we acknowledge that these models cannot replace the comprehensive qualifications, clinical experience, and human interactions essential to effective medical practice. Our findings should not be interpreted as an endorsement for using LLMs as substitutes for licensed medical professionals.
We recognize the potential risks associated with deploying LLMs in healthcare settings, including the dissemination of inaccurate or misleading information. To mitigate these risks, we emphasize that any application of LLMs in medical contexts should involve oversight by qualified healthcare practitioners and adhere strictly to ethical guidelines and regulatory standards.
Appendix A Detailed specialty performance on PES

### v2

@@ 5 Cross-lingual knowledge transfer
To compare the performance of various LLMs on Polish and English versions of the same datasets, we restrict the LEK and LDEK datasets to identical subsets. The English questions are translations of the original Polish questions, provided by human experts from the Medical Examination Center. Both versions are equivalent, meaning they convey the same medical content, structure, and intent, ensuring that the translated questions accurately reflect the original ones without altering their meaning or complexity. The analysis results, similar to the previous one, are presented in Tables 4 and 5. As shown, all medical models, except for OpenMeditron/Meditron3-70B, perform better on the English versions of the datasets. This may be due to these models being fine-tuned on English medical corpora. General-purpose multilingual models perform better on the English versions of the exams as well. This result is anticipated since these models are trained on corpora containing significantly more English than Polish. While these models are proficient in Polish, their performance on the tests remains lower in Polish than in English. The difference can be considerable; for example, meta-llama-Meta-Llama-3.1-8B-Instruct passed only one LEK exam in Polish but passed all 13 when translated into English. However, as model quality improves, the performance gap between languages narrows. For instance, with meta-llama-Meta-Llama-3.1-8B-Instruct, the accuracy difference between Polish LEK (51.25%) and English LEK (64.69%) is 13.44 percentage points (or a 26% relative change). In contrast, with meta-llama-Meta-Llama-3.1-70B-Instruct, the difference is only 1.66 percentage points (80.94% for Polish LEK vs. 82.60% for English LEK, or a 2% relative change).
For GPT-4o-mini, which generally performs well, the results in English are only slightly better than in Polish. Interestingly, for GPT-4o, performance is actually higher on the Polish version. The only Polish LLM, Bielik, performs better on Polish LEK and slightly better on Polish LDEK, likely due to its fine-tuning from the multilingual model Mistral-7B-v0.2 specifically for Polish. This fine-tuning enables it to better capture the nuances of Polish text to other models with a similar number of parameters. However, the tested Bielik-v2.2-Instruct, with only 11B parameters, is outperformed by models with double or even larger parameter counts on Polish versions of the LEK and LDEK exams. The only exceptions to this trend are Mistral-Small-Instruct-2409 and Llama3-OpenBioLLM-70B.
Overall, our observations suggest that language transfer is more effective as the model’s general performance improves. Refer to Appendix F for detailed question-level analysis.
@@
meta/llama-Meta-Llama-3.1-8B-Instruct 51.25 64.69 41.06 47.71
meta/llama-Meta-Llama-3.1-70B-Instruct 80.94 82.60 61.75 63.17
@@
mistralai/Mistral-Small-Instruct-2409 51.72 63.70 40.90 43.47
@@
Table 4: The comparison of LLMs on Polish and English datasets, using the same LEK and LDEK exams, is represented as a percentage of correct answers.
@@
meta-llama-Meta-Llama-3.1-8B-Instruct 1/13 13/13 0/13 0/13
meta-llama-Meta-Llama-3.1-70B-Instruct 13/13 13/13 12/13 13/13
@@
mistralai-Mistral-Small-Instruct-2409 1/13 13/13 0/13 0/13
@@
Table 5: The comparison of LLMs on Polish and English datasets using the same LEK and LDEK exams is represented as a passed exams.
@@ 6 Comparison against human results
Meditron3-70B, Meta-Llama-3.1-70B-Instruct, Bielik-11B-v2.2-Instruct, and gpt-4o-2024-08-06 are selected as the top-performing models for the groups mentioned in Section 4, and compared against anonymized human results published on the CEM webpage from the last four LEK and LDEK sessions (Spring 2024, Autumn 2023, Spring 2023, Autumn 2022), covering 977 LEK and 984 LDEK questions. The exams were taken by 33,929 participants for LEK and 4,366 for LDEK, totaling 38,295 results from medical graduates and final-year students in Poland. While all selected models pass the chosen LEK exams, only Meta-Llama-3.1-70B-Instruct and gpt-4o-2024-08-06 score within the range defined by an average number of points $\pm$ standard deviation achieved by humans. Assuming a normal distribution of exam results, it could be concluded that these models perform as a typical medical student. Notably, for the spring 2024 LEK exam, Meditron3-70B also achieves an average-level result, while gpt-4o-2024-08-06 exceeds the average student score. For the LDEK exams, all models perform noticeably worse. Assuming a normal distribution of exam results, only gpt-4o-2024-08-06 maintains a performance level comparable to that of an average medical student, consistent with its LEK exam results. In contrast, Meditron3-70B and Bielik-11B-v2.2-Instruct perform poorly, failing all exams, while Meta-Llama-3.1-70B-Instruct score below the average but manage to pass each exam. These outcomes are summarized in Table 7.
@@
The same models are used to compare their performance with humans on the PES exams. More details about joining the PES medical questions and human results are provided in Appendix G. The best-performing model is gpt-4o-2024-08-06, which achieves results in above 60% of cases better than half of the test takers population and above 30% of cases is placed in the top 25% of scores. Notably, this model outperforms all examinees in a thoracic surgery exam. However, it is important to note that the examinee population for this particular exam is relatively small, consisting of only six participants. However, it is worth noting that even the best model achieves results worse than half of the test takers population in over 30% of specializations. For the Audiology & phoniatrics specialization, the model underperformes compared to all examinees. However, the test takers population for that particular case was relatively small, consisting of only nine participants. The second-best model, Meta-Llama-3.1-70B-Instruct, delivers significantly worse performance compared to the best model. Only slightly above 11% of its results across specializations are above the population median, while in over 30% of medical specializations, its performance is above the 25th percentile and below the 50th percentile. The remaining models, Meditron3-70B and Bielik-11B-v2.2-Instruct, perform extremely poorly, with most of their results falling below the 25th percentile or even below the lowest scores of the entire test takers population. The human results and additional explanations for Table 6 are provided in Appendix G, where whiskers indicate the minimum and maximum human scores rather than the inter-quartile range.
@@
OpenMeditron/Meditron3-70B $Y<min(X)$ 17 25.00%
$Y\in[min(X),p_{25})$ 47 69.12%
$Y\in[p_{25},p_{50})$ 2 2.94%
$Y\in[p_{50},p_{75})$ 2 2.94%
$Y\in[p_{75},max(X))$ 0 0%
$Y\geq max(X)$ 0 0%
meta-llama/Meta-Llama-3.1-70B-Instruct $Y<min(X)$ 5 7.35%
$Y\in[min(X),p_{25})$ 33 48.53%
$Y\in[p_{25},p_{50})$ 22 32.35%
$Y\in[p_{50},p_{75})$ 6 8.83%
$Y\in[p_{75},max(X))$ 2 2.94%
$Y\geq max(X)$ 0 0%
speakleash/Bielik-11B-v2.2-Instruct $Y<min(X)$ 22 33.82%
$Y\in[min(X),p_{25})$ 44 64.71%
$Y\in[p_{25},p_{50})$ 1 1.47%
$Y\in[p_{50},p_{75})$ 0 0%
$Y\in[p_{75},max(X)])$ 0 0%
$Y\geq max(X)$ 0 0%
gpt-4o-2024-08-06 $Y<min(X)$ 1 1.47%
$Y\in[min(X),p_{25})$ 9 13.24%
$Y\in[p_{25},p_{50})$ 15 22.06%
$Y\in[p_{50},p_{75})$ 20 29.41%
$Y\in[p_{75},max(X))$ 22 32.35%
$Y\geq max(X)$ 1 1.47%
Table 6: Aggregated PES exam results categorizing model $Y$ performance relative to the test takers population $X$ across various percentiles, from scores below all examinees ($Y<min(X)$) to scores compared to or exceeding the best human results ($Y\geq max(X)$). Additional explanations are available in Appendix G.
@@
with standard deviation $\pm 19.79$ $\pm 18.38$ $\pm 18.66$ $\pm 16.95$
(a) LEK
@@
with standard deviation $\pm 26.08$ $\pm 19.08$ $\pm 21.13$ $\pm 23.52$
(b) LDEK
Table 7: Comparison of top-performing LLMs and average human results, including standard deviation, across selected LEK and LDEK exams. Red represents values below the passing threshold of 112 points, orange highlights scores below average minus one standard deviation, green indicates scores above average plus one standard deviation, and black represents scores within one standard deviation of the average.
7 Conclusion
In this paper, we propose a new benchmark for analyzing the performance of large language models in answering questions pertaining to the domain of medical knowledge. In contrast to the majority of previous medical datasets that collect examination questions in English, our dataset is derived from data of Polish origin. We show that general-purpose LLMs, trained on internet-scale datasets with extensive computational resources, outperform medical-specific models and that using a general-purpose model fine-tuned specifically for the Polish language is justified only if models of a similar size are considered.
LLMs performance varies across different medical exams and languages. Most models are able to pass the LEK exam, but many struggle with the LDEK exam. In the case of the PES exams for various medical specializations, the performance is even lower, with only gpt-4o-2024-08-06 maintaining satisfactory results. However, even the top-performing model scores lower than at least half of the test takers in over 30% of specializations. This highlights the need for thorough verification before implementing LLMs in the medical domain, as the results are not consistently reliable across all medical specialties.
The parallel sub-corpus composed of examination questions in Polish aligned with their English counterparts is a distinguished feature of the presented benchmark which allows us to investigate the cross-lingual transfer of medical knowledge in LLMs. Our findings show that models perform significantly better on English questions and that as the size of the model increases, performance improves, and the gap between languages narrows, an expected but difficult-to-measure result without an appropriate benchmark.
@@ 7 Conclusion
While LLMs have demonstrated impressive performance on Polish medical multiple-choice exams, this achievement represents only a narrow facet of medical expertise. Becoming a licensed physician in Poland requires extensive training, rigorous coursework, and hands-on experience with practical medical procedures—far beyond what written exams can assess. Clinical practice necessitates analyzing diverse information and solving complex problems with multiple possible solutions. Physicians must determine what data is needed, obtain it through patient interviews, physical examinations, diagnostic tests, and consultations—all heavily reliant on direct human interaction that AI models cannot replicate. Moreover, the exams are multiple-choice, and real-world work is not narrowed to a few possible options. Therefore, despite strong exam results, LLMs cannot currently substitute the comprehensive qualifications and essential human interactions integral to effective medical care. However, this study demonstrates that LLMs may serve as valuable tools for medical practitioners, a potential use case previously suggested by other researchers Ullah et al. (2024); Park et al. (2024); Clark and Bailey (2024); Liu et al. (2023); Lee et al. (2023).
@@
The presented research and provided benchmark are primarily designed to evaluate the performance of LLM models in medical question-answering tasks in the Polish language, serving as a substitute for the MedQA benchmark. To prevent models from being trained on the benchmark data, the training dataset is not provided. The LEK and LDEK exams have a similar number of examples in the final dataset, but the PES exam questions make up the majority. While the English questions for the LEK and LDEK exams are in the minority, there are no PES questions in English. This benchmark is not intended for other NLP tasks such as text classification, named entity recognition, sentiment analysis, or other related problems.
The GPT-4o model accessed via the OpenAI API was used for our analysis. The web search option was not enabled, ensuring that the model did not actively search the Internet for answers to the medical questions asked.
The dataset described in this paper was collected from an examination center’s webpage, where the questions are freely available. These exams can be used, for example, by medical students preparing for their assessments. There is a potential risk for these exams being included in the training datasets of evaluated LLMs. Therefore, the evaluation results presented in this paper must be treated with the same degree of caution as the results determined with the use of any other publicly available dataset such as MMLU Hendrycks et al. (2020) or scores reported on leaderboards that aggregate results determined for publicly available datasets, such as Open LLM Leaderboard Fourrier et al. (2024). However, taking into consideration that our dataset originates from a highly authoritative source, creating a dataset of comparable size and quality from the ground up would be prohibitively difficult.
Appendix A Polish medical exams detailed description
https://www.cem.edu.pl/ldek_info.php
Both LEK and LDEK are four-hour exams conducted twice a year. Each exam consists of 200 multiple-choice questions with five possible answers, of which only one is correct. The questions cover a wide range of medical or dental disciplines. The distribution of questions from various fields is presented in Tables 8 and 9. To pass, a candidate must correctly answer at least 56% of the questions. Physicians and dentists can retake these exams multiple times, even after passing, if they are dissatisfied with their score. [^4] https://www.cem.edu.pl/lek_info.php
https://www.cem.edu.pl/ldek_info.php
A controversial rule (https://pulsmedycyny.pl/kadry/lekarze/samorzad-lekarski-postuluje-pilna- zmiane-bazy-pytan-w-lek-i-ldek/) has been introduced in 2022, stipulating that 70% of the exam questions come from a publicly available database, which includes 2,870 questions for LEK and 3,198 for LDEK. After these changes, the average exam scores and the percentage of passing candidates increased significantly. [^5] https://www.cem.edu.pl/lep_s_h.php
https://www.cem.edu.pl/ldep_s_h.php
Discipline Questions
Internal medicine* 39
Pediatry* 29
Surgery* 27
Obstetrics and gynecology* 26
Psychiatry 14
Family medicine* 20
Emergency medicine and intensive care 20
Bioethics and medical law 10
Medical certification 7
Public health 8
Discipline Questions
Conservative dentistry* 46
Pediatric dentistry* 29
Oral surgery* 25
Prosthetic dentistry 25
Periodontology* 20
Orthodontics* 20
Emergency medicine 10
Bioethics and medical law 10
Medical certification 7
Public health 8
Appendix B Example exam questions
B.1 LEK
Exam: 2022 Spring
Question id: 77
Przepuklina u starszego mczyzny
z chorob obturacyjn puc uwypuklajca
si na zewntrz jamy brzusznej przez
powi poprzeczn stanowic tyln
cian kanau pachwinowego w miejscu
ograniczonym od gry przez cigno
czce, od dou przez wizado
pachwinowe, a bocznie przez naczynia
nabrzuszne dolne - jest rozpoznawana
jako:
A. przepuklina pachwinowa skona.
B. przepuklina mosznowa.
C. przepuklina pachwinowa prosta.
D. przepuklina udowa.
E. przepuklina Spigela.
Correct answer: C.
B.2 LEK (en)
This sample is a translation of the above question (LEK) provided by the examination center.
Exam: 2022 Spring
Question id: 77
An elderly male patient with obturative
lung disease was diagnosed with hernia.
It was protruding from the abdominal
cavity through the transverse fascia
which forms the posterior wall of the
inguinal canal, at the site bordering
the conjoint tendon at the top, the
inguinal ligament at the bottom, and
laterally, through inferior  epigastric
vessels. The hernia in such location is
known as:
A. oblique inguinal hernia.
B. scrotal hernia.
C. direct inguinal hernia.
D. femoral hernia.
E. spigelian hernia.
Correct answer: C.
B.3 LDEK
Exam: 2022 Spring
Question id: 77
Jednostronny wyciek z nosa posokowatej
treci z domieszk krwi, rozchwianie
zbw grnych, zawienie, wytrzeszcz
gaki ocznej, a niekiedy ble i
mrowienie policzka mog by wczesnym
objawem:
A. pseudotorbieli zatoki szczkowej.
B. raka zatoki szczkowej.
C. raka policzka.
D. przewlekego zapalenia zatoki szczkowej.
E. ostrego zapalenia zatoki szczkowej.
Correct answer: B.
B.4 LDEK (en)
This sample is a translation of the above question (LDEK) provided by the examination center.
Exam: 2022 Spring
Question id: 77
Unilateral ichorous discharge from the
nose with a blend of blood, gomphiasis
of the upper teeth, lacrimation,
exopathalmos, and sometimes pain and
tingling sensation in the cheek, might
be an early symptom of:
A. pseudocyst of the maxillary sinus.
B. cancer of the maxillary sinus.
C. buccal cancer.
D. chronic maxillary sinusitis.
E. acute maxillary sinusitis.
Correct answer: B.
B.5 PES
Exam: 2019 Autumn
Question id: 68
Specialty: Family medicine
 Kliniczne cechy sepsy u dzieci to:
1) gorczka;
2) leukocytoza;
3) leukopenia;
4) tachykardia bez innej przyczyny;
5) tachypnoe bez innej przyczyny.
Correct answer: E.
Appendix C Prompts
C.1 Prompt in Polish
Twoje zadanie to udzielenie odpowiedzi
na test medyczny dla lekarzy. Spord
wszystkich odpowiedzi A,B,C,D,E wybierz
tylko jedn. Jeeli nie jeste pewien,
wybierz najbardziej prawdopodobn.
Odpowiedz w sposb:
Prawidowa odpowied to B.
C.2 Prompt in English
Your task is to answer a medical test
for doctors. From all the options
A, B, C, D, E, choose only one. If you’re
unsure, select the most probable one.
Respond in the following manner:
The correct answer is B.
Appendix D Specialty performance on PES
Among the 72 unique PES specialties, certain areas of medicine consistently challenge the majority of tested models, while others frequently rank among the highest-scoring categories based on model accuracy. By identifying the top five highest and lowest-scored categories, we gain insights into specific domains where models excel or struggle, highlighting their potential limitations in these fields.